Glenn budgets 1/6 of his weekly allowance for entertainment. Glenn’s entertainment bill is $29. What is his weekly allowance?

Sheffield Corporation sells sets of pots ($14) and dishes ($12) at a local charity. On Labor Day weekend. Sheffield’s total sales were $1,080. People bought three times as many pots as dishes. How many dishes did Sheffield sell?
Brian Smith sells sets of movie posters ($12) and baseball cards ($8) at a local flea market. On last Sunday, Brian’s total sales were $840. People bought four times as many movie posters as baseball cards. How much did Brian make on the movie posters?

Answer by ikleyn(52781)

.

            The GOLDEN  RULE  of this forum is  ONE  and  ONLY  ONE  problem per post.

            Therefore,  I will solve only one problem in my post.

            PLEASE  DO  NOT  VIOLATE  THE  RULE.


    Sheffield Corporation sells sets of pots ($14) and dishes ($12) at a local charity.  On Labor Day weekend,  
    Sheffield’s total sales were $1,080.  People bought three times as many pots as dishes.  How many dishes did Sheffield sell? 


Solution

Let x be the number of dishes.

Then the number of pots is  3x,  according to the condition.


Thus your "money" equation is


    dishes + pots    = total revenue,      or


    12*x   + 14*(3x) = 1080   dollars.


    12x + 42x = 1080

    56x = 1080

    x = 1080/54 = 20.


ANSWER.  20 dishes.

Solved.

--------------

The next problem regarding "Brian Smith selling sets of movie posters and baseball cards",  is a TWIN to this solved problem.

Therefore,  you can solve  "the Brian Smith"  problem  BY  THE  SAME  WAY,  using my solution above as a  TEMPLATE.


Good luck and happy solving !


Have a nice day !